Hi, Could you help me with this proof?
1.
2.
3. AC=DE
4.GH=GH
5. AC + GH = DE + GH
I need to know the reason for number five.
these are the choices:
A. Symmetric Property of Equality
B. Substitution Property of Equality
C. Addition Property of Equality
D. Division Property of Equality
I am really confused, as my tutor is out of town. if you could tell me which one and why it would really help. thanks in advance.

Answer by scott8148(3755)

the reason is choice C. Addition Property of Equality

you have two equal quantities which are ADDED to two other equal quantities ... the result must be equal quantities

4=4 ... 7=7 ... 4+7=4+7 ... 11=11